# Torque-based Graph Surgery: Enhancing Graph Neural Networks with Hierarchical Rewiring

### v1

@@ Abstract
Graph Neural Networks (GNNs) have emerged as powerful tools for learning from graph-structured data, leveraging message passing to diffuse information and update node representations. However, most efforts have suggested that native interactions encoded in the graph may not be friendly for this process, motivating the development of graph rewiring methods. In this work, we propose a torque-driven hierarchical rewiring strategy, inspired by the notion of torque in classical mechanics, dynamically modulating message passing to improve representation learning in heterophilous graphs and enhance robustness against noisy graphs. Specifically, we define an interference-aware torque metric that integrates structural distance and energy scores to quantify the perturbation induced by edges, thereby encouraging each node to aggregate information from its nearest low-energy neighbors. We use the metric to hierarchically reconfigure each layer’s receptive field by judiciously pruning high-torque edges and adding low-torque links, suppressing propagation noise and boosting pertinent signals. Extensive evaluations on benchmark datasets show that our approach surpasses state-of-the-art methods on both heterophilous and homophilous graphs, and maintains high accuracy on noisy graph.
@@ 1 Introduction
In practice, however, raw graphs frequently harbour spurious or missing links arising from noise or sampling artefacts, compromising their effectiveness as substrates for message propagation. In response, recent work has devised diverse graph rewiring strategies that selectively remove and add edges to optimize message passing and boost predictive accuracy [11, 12, 13, 14]. Such dynamic topology adjustment is crucial not only for mitigating noise but also for addressing heterophily, where nodes with dissimilar labels or features tend to be connected [15, 16, 17]. In such scenarios, homophily-based GNNs can be misled by abundant heterophilous connections, yielding entangled representations and degraded classification accuracy.
To explore the factors that influence message propagation for guiding effective rewiring, we simulate attacks by injecting adversarial edges into raw graphs and visualize the resulting structural perturbations. As shown in Figure 1(a)-(b), adversarial edges in the homophilous Cora span node pairs with larger feature distances than native connections. A similar trend is observed in Wisconsin that inherently includes many heterophilous edges. To further uncover the attacker’s intention, we estimate node importance via the PageRank algorithm [18, 19] and derive an edge-level PageRank score by aggregating scores of each edge’s endpoints. As illustrated in Figure 1(c)-(d), adversarial edges exhibit two distinct patterns: in Cora, they preferentially connect high-PageRank nodes; whereas in Wisconsin, they appear predominantly among low-PageRank nodes, suggesting that PageRank cannot fully reveal the essence of adversarial edges. Moreover, real-world graphs inherently include spurious links even without external attacks, rendering PageRank insufficient to accurately assess the disruptive potential of edges. This motivates a key question: can we jointly leverage structural distance and a more expressive metric of edge perturbation to achieve fine-grained graph refinement?
@@
Figure 1: Density distributions of distances and PageRank scores for normal vs. adversarial edges on Cora and Wisconsin, where edge PageRank is the sum of endpoint scores in Cora and the absolute difference in Wisconsin.
@@
We draw inspiration from the concept of Torque in classical mechanics, mathematically defined as the cross product of a lever arm (the position vector from the axis of rotation to the point of force application) and a force. Recently, torque has found utility in biology [20, 21, 22] and spintronics [23, 24]. Heuristically, we extend this concept to graphs by treating the distance between nodes as the lever arm and the noise exerted by a neighboring node on a central node as the force. Their product yields a graph torque, which measures an edge’s disruptive intensity: higher torque flags greater interference. To our knowledge, this is the first work to integrate a physics-inspired torque into graph rewiring, enabling an interference-aware message passing.
@@
Specifically, we devise a Torque-driven Hierarchical Rewiring strategy (THR) for GNNs (TorqueGNN), which dynamically refines message passing to excel in both heterophilous and noisy graph. In THR, each edge is endowed with two learnable attributes: its Euclidean distance as the “lever arm”, which accentuates long-range or heterophilous links, and an energy score as the “force”, which is computed by its endpoints reflecting high values to noisy or noise-sensitive nodes and low values to benign ones. Their product defines a torque to quantify the interference strength of edges, where larger torques flag less reliable connections. Leveraging this metric, TorqueGNN hierarchically reconfigures each layer’s receptive field via automatically removing undesirable edges that result in performance degrading and introducing low-torque significant connections, thus effecting interference-resistant and importance-aware propagation. This rewiring is performed end-to-end, where message passing operates on the continuously updated graph, while the evolving node representations enhance torque computation.
@@
Contributions: 1) To our knowledge, we are the first effort to apply the torque in physics into graph rewiring, culminating in TorqueGNN that enhances GNNs resilience to both heterophily and noise. 2) We devise a hierarchical rewiring strategy that adaptively infers each layer’s receptive field through automatically pruning undesirable connections and adding significant edges. 3) Comprehensive experiments indicate that TorqueGNN outperforms advanced GNNs across standard benchmarks and maintains robustness in noisy graphs, with THR surpassing superior rewiring strategies.
@@ 2 Related Work
Standard message passing in GNNs, which aggregates information from local neighbourhoods, struggles to capture long-range dependencies. A common remedy is to stack multiple layers to expand the receptive field [25, 26, 27], but this approach frequently encounters fundamental limitations such as over-smoothing and over-squashing. To address these bottlenecks, graph rewiring techniques have recently emerged as a powerful strategy to restructure connectivity and enhance information flow. For example, Expander GNNs and ExPhormer perform graph rewiring by merging multi-hop neighbourhoods or injecting virtual nodes [28, 29, 30]. [31] adds edges based on spectral expansion to mitigate over-smoothing and over-squashing, while degree-preserving local edge-flip algorithms are developed [32]. [33, 34] analyze the root causes of over-squashing, demonstrating that both spatial and spectral rewiring can effectively counteract this bottleneck. Despite these advances, real-world graphs often remain inherently noisy, limiting the reliability of message passing.
Raw graph structures are frequently suboptimal substrates for message passing, as noisy edges and perturbed features can induce both spurious and missing links, thereby limiting the expressive power of GNNs [33, 35]. To mitigate the effect of such undesirable connections, [36] simultaneously perform graph rewiring and feature denoising to boost classification accuracy. Moreover, [36] highlights the challenge posed by heterophilous edges, wherein the aggregation of dissimilar node signals can lead to entangled representations and misclassifications. There are many methods specifically devised for heterophily. For example, [37, 38, 39] employ signed message propagation, assigning positive weights to homophilous links and negative weights to heterophilous ones. This enables differentiated updates on the heterophilous graphs, thereby amplifying similarity among homophilous nodes while suppressing similarity among heterophilous ones. However, [40] has shown that, although single-hop signed adjacency matrix aids in distinguishing features of different classes, the multi-hop propagation matrix introduced to expand the receptive field often degrades performance.
We are inspired by the torque in physics to design a new rewiring mechanism that hierarchically eliminates undesirable connections and incorporates task-relevant edges. By dynamically reshaping the receptive field during training, our method enhances the discriminative power of GNNs in noisy and heterophilous graphs.
@@ 3.2 Message Passing
[…]
3.3 Energy-based Model
Energy-based model (EBM) [42] aims to establish an energy function $E(\mathbf{x},y):\mathbb{R}^{d}\rightarrow\mathbb{R}$, which maps an input instance $\mathbf{x}$ with a given label $y$ to a scalar value termed energy. Consequently, the Boltzmann distribution is expressed as
$p(y|\mathbf{x})=\frac{\text{exp}\big{(}-E(\mathbf{x},y)\big{)}}{\sum_{y^{\prime}}\text{exp}\big{(}-E(\mathbf{x},y^{\prime})\big{)}}=\frac{\text{exp}\big{(}-E(\mathbf{x},y)\big{)}}{\text{exp}\big{(}-E(\mathbf{x})\big{)}},$ (2)
where the denominator $\sum_{y^{\prime}}\text{exp}\big{(}-E(\mathbf{x},y^{\prime})\big{)}$ is the partition function. [43, 44] connect neural classifiers $f(\cdot)$, which map an input to a $c$-dimensional logit vector, with energy-based models by setting the energy as the negative logit value of the classifier $E(\mathbf{x},y)=-f(\mathbf{x})[y]$, where $f(\mathbf{x})[y]$ denotes the non-normalized logit corresponding to class $y$. Then, the energy of the input instance marginalizing over all possible labels becomes
$E(\mathbf{x})=-\text{log}\sum_{y^{\prime}}\text{exp}\big{(}-E(\mathbf{x},y^{\prime})\big{)}=-\text{log}\sum_{y^{\prime}}\text{exp}\big{(}f(\mathbf{x})[y^{\prime}]\big{)}.$ (3)
@@ 4 Methodology
In this section, we propose a novel graph rewiring strategy that unfolds in three key stages: (i) computing edge torques, (ii) rewiring propagation matrix, and (iii) adjusting message passing. The full algorithmic pseudocode is provided in Appendix A, while a detailed complexity analysis of TorqueGNN is included in Appendix B.
@@ 4.1 Derive Graph Torque
where $\mathbf{r}$ denotes the lever arm vector connecting the point of application to the axis of rotation, $\mathbf{F}$ indicates the force vector, and $\theta$ is the angle between them. The magnitude of the torque directly governs an object’s tendency to rotate under the applied force. In GNNs, the systematic exploration of node interactions mirrors the lever arm–force relationship underlying torque in classical mechanics. Specifically, we can transpose torque to graphs by treating the distance $D_{\langle i,j\rangle}$ between a central node $v_{i}$ and its neighbor $v_{j}$ as a scalar “lever arm”, and the neighbor’s influence level $E_{\langle i,j\rangle}$ as a scalar “disturbing force” on $v_{i}$. In graphs, the notion of a disturbing force is abstract and lacks a well-defined direction, as message passing in GNNs does not involve physical "rotation" in the classical mechanical sense. Therefore, rather than computing the full vector torque via a cross product, we adopt a scalar approximation based on the product of distance and energy. Mathematically, for the edge $k$ connecting nodes $v_{i}$ and $v_{j}$, we define its torque value as follow,
$T_{e_{k}}=D_{\langle i,j\rangle}\times E_{\langle i,j\rangle},$ (5)
which measures the disturbance strength that the message carried by edge $e_{k}$ imposes on node $v_{i}$. Obviously, larger distances or higher disturbance both amplify torque, while edges maximizing both factors yield the greatest torque that represents the highest priority for graph rewiring. Crucially, the definition of $D$ and, in particular, $E$ in Eq. (5), remains a key challenge for accurate rewiring. Here, we employ an EBM to quantify node interference, and subsequently delineate the construction of both $D$ and $E$.
Metric 1: Distance. To mitigate the effect of noise in raw graphs and features, we estimate node distances using optimized representations. Specifically, for a central node $v_{i}$ and its neighbor $v_{j}$, we have the following formulation
$D_{\langle i,j\rangle}=\|\mathbf{h}_{i}-\mathbf{h}_{j}\|_{2},$ (6)
@@
Metric 2: Energy. While graph-based message passing can markedly enhance node representations, it also risks amplifying structural and feature noise. To mitigate this, we posit nodes that persistently absorb noisy signals exhibit elevated interference strength relative to benign nodes. EBM, which assigns scalar energy values to inputs and has demonstrated remarkable robustness against adversarial perturbations [45, 46], provide a foundation for this purpose. Therefore, we introduce an EBM-derived energy score as another indicator of edge reliability, as shown in Eq. (3), which assigns low energy to benign samples and high energy to corrupted ones. In our framework, an energy-based GNN jointly performs node classification and energy estimation. For a network parameterized by $\mathbf{\Theta}$, the energy of node $v_{i}$ is defined as $E_{\mathbf{\Theta}}(\mathbf{x}_{i})=-\text{log}\sum_{y}\text{exp}(\mathbf{h}_{i}[y]\big{)}$. Here, $\mathbf{h}_{i}$ denotes the logit vector for node $v_{i}$. We then derive an edge-level energy metric by combining the energy scores of node pairs,
$E_{\langle i,j\rangle}=E_{\mathbf{\Theta}}(\mathbf{x}_{i})\times E_{\mathbf{\Theta}}(\mathbf{x}_{j}).$ (7)
@@ 4.2 Adjust Message Passing
Edge-removal High-order Rewiring. Herein, we propose an automated threshold learning mechanism that identifies the optimal number of edges to prune by pinpointing the largest successive torque gap. Specifically, we first rank all $K$ edges in descending order of their torque values to form a torque-sorted list (TSL), denoting its $k$-th entry entry as $\widetilde{e}_{k}$ with torque $\widetilde{T}_{e_{k}}$, so that $\widetilde{T}_{e_{1}}\geq\widetilde{T}_{e_{2}}\geq\cdots\geq\widetilde{T}_{e_{K}}$. We then calculate the torque gap between two consecutive links by
@@
where $\delta$ is a small constant to prevent division by zero, and the weight $\mu_{k}$ reflects the proportion of anomalous edges, those whose distance $D$, energy $E$ and torque $T$ all exceed their respective means, that are captured within the top $k$ torque-ranked set, emphasizing the boundary between desirable and undesirable connections. The computation formula of $\mu_{k}$ is defined as
@@
where $\bar{D},\bar{E},\bar{T}$ denote the mean values of distance, energy and torque, respectively, computed over all $K$ edges. The set $High\_e$ comprises edges exhibiting above-average values across all three metrics, while $Top\_k$ contains the top $k$ connections in TSL. According to (8), we can identify the optimal cutoff by locating the largest torque gap $\mathcal{K}=\arg\max\limits_{0\leq k\leq K-1}G_{k,k+1}$, which separates the edge set into two groups: undesirable connections ($\widetilde{e}_{1},\cdots,\widetilde{e}_{\mathcal{K}}$) and desirable connections ($\widetilde{e}_{\mathcal{K}+1},\cdots,\widetilde{e}_{K}$).
In practice, multi-layer GNNs, such as APPNP [47], are widely adopted to enlarge the receptive field of graph convolutions. Building on the pairwise torque gap formulation introduced above, we extend this mechanism across multiple propagation layers. In specific, for each order $l$, we construct a dedicated propagation matrix that enables selective filtering of undesirable high-order interactions. Let $\mathcal{A}^{(l)}$ denote the refined propagation matrix used in $l$-th layer and $\mathcal{A}^{(0)}=\mathbf{A}$. We have current neighbors $v_{j}$ of a given node $v_{i}$ based on $\mathcal{A}^{(l)}$, and recompute two key properties as follows
$D_{\langle i,j\rangle}^{(l+1)}=\|\mathbf{h}_{i}^{(l)}-\mathbf{h}_{j}^{(l)}\|_{2},\quad E_{\langle i,j\rangle}^{(l+1)}=E_{\mathbf{\Theta}}^{(l)}(\mathbf{x}_{i})\times E_{\mathbf{\Theta}}^{(l)}(\mathbf{x}_{j}),$ (10)
[…]
@@
[…]

### v3

# GraphTorque:: Torque-Driven Rewiring Graph Neural Network

@@ Abstract
Graph Neural Networks (GNNs) have emerged as powerful tools for learning from graph-structured data, leveraging message passing to diffuse information and update node representations. However, most efforts have suggested that native interactions encoded in the graph may not be friendly for this process, motivating the development of graph rewiring methods. In this work, we propose a torque-driven hierarchical rewiring strategy, inspired by the notion of torque in classical mechanics, dynamically modulating message passing to improve representation learning in heterophilous and homophilous graphs. Specifically, we define the torque by treating the feature distance as a “lever arm vector” and the neighbor feature as a “force vector” weighted by the homophily disparity between nodes. We use the metric to hierarchically reconfigure each layer’s receptive field by judiciously pruning high-torque edges and adding low-torque links, suppressing the impact of irrelevant information and boosting pertinent signals during message passing. Extensive evaluations on benchmark datasets show that the proposed approach surpasses state-of-the-art rewiring methods on both heterophilous and homophilous graphs.
@@ 1 Introduction
In practice, however, raw graphs frequently harbour spurious or missing links arising from noise or sampling artefacts, compromising their effectiveness as substrates for message propagation. In response, recent work has devised diverse graph rewiring strategies that selectively remove and add edges to optimize message passing and boost predictive accuracy [11, 12, 13, 14, 15]. Such dynamic topology adjustment is crucial not only for mitigating spurious connections but also for addressing heterophily, where nodes with dissimilar labels or features tend to be connected [16, 17, 18]. In such scenarios, homophily‑based GNNs can be misled by abundant heterophilous connections, yielding entangled representations and degraded classification accuracy.
One of the core challenges in graph rewiring is quantifying the impact of edges on message passing. A key factor in this process is the similarity between node pairs, often measured using the Euclidean distance, a commonly used metric for assessing similarity. In general, the larger the distance between nodes, the weaker their interaction strength, and the less useful information can be transmitted, as supported by previous studies that employed node similarity as a proxy for edge weights [19, 20]. To intuitively observe this, we simulate adversarial attacks by injecting adversarial edges into raw graphs and visualize the distance distribution of the edges, enabling us to examine whether adversarial and original edges exhibit distinct distributional patterns. As shown in Fig. 1(a)–(d), the distribution trends in both homophilous datasets (Cora and PubMed) and heterophilous datasets ( Wisconsin and Texas ) consistently indicate that adversarial edges (in red) tend to connect node pairs with larger feature distances. These observations suggest that adversarial attacks preferentially create long-range links so that they disrupt message passing at their target nodes. Furthermore, we observe that normal edges in heterophilous datasets also exhibit a distribution skewed toward larger distances, more pronounced than in homophilous datasets. This is because heterophilous graphs contain a much higher proportion of heterophilous edges, which typically span node pairs with low similarity (i.e., large distances). Given that a minority of long-range neighbors can convey crucial information while nearby neighbors may propagate misleading signals, the feature quality of neighboring nodes should be another key factor in assessing edge significance.
@@
Figure 1: Density distributions of distances for normal vs. adversarial edges on homophilous graphs (a) Cora and (b) Pubmed and heterophilous graphs (c) Wisconsin and (d) Texas.
@@
This brings to mind the concept of Torque in classical mechanics, which is mathematically defined as the cross product of a lever arm (the position vector from the axis of rotation to the point of force application) and a force. Recently, torque has found applications in fields such as biology [21, 22, 23] and spintronics [24, 25]. Heuristically, we extend this concept to graphs by treating the distance vector between nodes as the lever arm and the feature vector of a neighboring node as the force. Their product yields a graph torque, which measures an edge’s negative impact: higher torque flags greater interference. To our knowledge, this is the first work to integrate a physics-inspired torque into graph rewiring, enabling an interference-aware message passing.
@@
Specifically, we devise a Torque‑driven Hierarchical Rewiring strategy (THR) for GNNs, which dynamically refines message passing to excel in both homophilous and heterophilous graph. In THR, each edge is assigned a torque value that quantifies its interference strength, with larger torques indicating less reliable connections. Sepcifically, we define torque by treating the difference between node representations as a “lever arm vector”, which emphasizes long-range or heterophilous links. Meanwhile, the neighbor feature is treated as a “force vector” weighted by the disparity in the homophily ratios between node pairs. This disparity captures the difference in their local label homophily, which has been theoretically shown to jointly influence the expressive power of GNNs together with feature distance. Leveraging this torque value, THR hierarchically reconfigures each layer’s receptive field via automatically removing undesirable edges that degrade performance and introducing low-torque significant connections, thus effecting interference-resistant and importance-aware propagation. This rewiring is performed end-to-end, where message passing operates on the continuously updated graph, while the evolving node representations enhance torque computation.
@@
Contributions: 1) To the best of our knowledge, we are the first to apply the concept of torque from physics to graph rewiring, resulting in THR, which enhances GNNs’ resilience to both homophily and heterophily. 2) We propose a hierarchical rewiring strategy that adaptively determines each layer’s receptive field by automatically pruning undesirable connections and adding significant edges. 3) Comprehensive experiments indicate that THR improves the performance of various GNNs and outperforms existing state-of-the-art rewiring strategies.
@@ 2 Related Work
Standard message passing in GNNs, which aggregates information from local neighbourhoods, struggles to capture long-range dependencies. A common remedy is to stack multiple layers to expand the receptive field [26, 27, 28], but this approach frequently encounters fundamental limitations such as over‑smoothing and over‑squashing. To address these bottlenecks, graph rewiring techniques have recently emerged as a powerful strategy to restructure connectivity and enhance information flow. For example, Expander GNNs and ExPhormer perform graph rewiring by merging multi‑hop neighbourhoods or injecting virtual nodes [29, 30, 31]. [32] adds edges based on spectral expansion to mitigate over-smoothing and over-squashing, while degree-preserving local edge-flip algorithms are developed [33]. [34, 35] analyze the root causes of over-squashing, demonstrating that both spatial and spectral rewiring can effectively counteract this bottleneck.
Moreover, [36] highlights the challenge posed by heterophilous edges, wherein the aggregation of dissimilar node signals can lead to entangled representations and misclassifications. To mitigate the effect of such undesirable connections, [14] compares the neighborhood feature distribution and neighborhood label distribution between node pairs to pruning heterophilous edges and adding homophilous edges. [37, 38, 39] employ signed message propagation, assigning positive weights to homophilous links and negative weights to heterophilous ones. This enables differentiated updates on the heterophilous graphs, thereby amplifying similarity among homophilous nodes while suppressing similarity among heterophilous ones. However, [40] has shown that, although single‑hop signed adjacency matrix aids in distinguishing features of different classes, the multi‐hop propagation matrix introduced to expand the receptive field often degrades performance.
We are inspired by the torque in physics to design a new rewiring mechanism that hierarchically eliminates undesirable connections and incorporates task‑relevant edges. By dynamically reshaping the receptive field during training, our method enhances the discriminative power of GNNs on both homophilous and heterophilous graphs.
@@ 3.2 Message Passing
where $\mathbf{h}_{i}^{(0)}=\mathbf{x}_{i}$, and $\mathbf{h}_{i}^{(l+1)}\in\mathbb{R}^{m}$ is the representation of node $v_{i}$ in the $(l+1)$-th layer. $\text{Agg}(\cdot)$ computes the incoming message from a neighbor $v_{j}$, and $\text{Upd}(\cdot)$ updates the representation of node $v_{i}$. Rather than relying on the raw adjacency matrix $\mathbf{A}$, most GNNs adopt a modified propagation operator $\mathcal{A}$. For example, GAT [41] replaces each non-zero entry of $\mathbf{A}$ with a learned attention coefficient that depends on the representations of the corresponding node pair.
3.3 Node-level Homophily and Heterophily
For a set of nodes with labels, the homophily ratio of each node quantifies the tendency of the node to share the same label as its neighbors. Considering a node $v_{i}$ and its set of neighbors $\mathcal{N}_{i}$, the homophily ratio $h_{i}^{+}$ of $v_{i}$ is defined as: $h_{i}^{+}=\frac{|\{\mathbf{y}_{i}=\mathbf{y}_{j}|v_{j}\in\mathcal{N}_{i}\}|}{|\mathcal{N}_{i}|}$. The value of $h_{i}^{+}$ lies in the range $[0,1]$, where values closer to 1 indicate a higher degree of homophily (or lower heterophily), while values nearer to 0 signify the opposite. To quantify the homophily of the entire graph $\mathcal{G}$, we compute the average homophily across all nodes: $\mathcal{H}(\mathcal{G})=\frac{\sum_{i=1}^{N}h_{i}^{+}}{N}$.
@@ 4 Methodology
In this section, we propose a novel graph rewiring strategy that unfolds in three key stages: (i) computing edge torques, (ii) rewiring propagation matrix, and (iii) adjusting message passing. The full algorithmic pseudocode is provided in Appendix B.
@@ 4.1 Derive Graph Torque
where $\mathbf{r}$ denotes the displacement vector, $\mathbf{F}$ indicates the force vector, and $\theta$ is the angle between them. The magnitude of the torque directly governs an object’s tendency to rotate under the applied force. In GNNs, the systematic exploration of node interactions can mirror the lever arm–force relationship underlying torque in classical mechanics. Specifically, torque on a graph may be conceptualized by treating the displacement vector $\mathbf{D}_{\langle i,j\rangle}$ between a central node $v_{i}$ and its neighbor $v_{j}$ as the effective “lever arm”, while the neighbor’s features $\mathbf{x}_{j}$ act as the applied “force". However, the contribution of this force varies across different central nodes. Recent studies [42, 7] demonstrate that the generalization of GNNs is influenced by two key factors: the proximity of aggregated features and the disparity in homophily ratio, with smaller values yielding better generalization. Inspired by this, we introduce the homophily ratio disparity term $E_{\langle i,j\rangle}$ to modulate the effective force, thereby not only capturing the heterogeneous influences of neighboring nodes but also unifying these two factors within the torque formulation to reduce generalization error.
Mathematically, for an edge $k$ connecting nodes $v_{i}$ and $v_{j}$, we define the corresponding torque as follow,
$\mathbf{T}_{e_{k}}=\mathbf{D}_{\langle i,j\rangle}\times E_{\langle i,j\rangle}\mathbf{x}_{j}.$ (3)
Its magnitude, denoted $T_{e_{k}}$, measures the disturbance imposed by the message passing along edge $e_{k}$ on node $v_{i}$. The value naturally increases with larger distances or higher structural disparity, with edges maximizing both factors yielding the greatest torque value that represents the highest priority for graph rewiring. A central goal is therefore to provide a principled definition of the displacement vector $\mathbf{D}_{\langle i,j\rangle}$ and the homophily ratio disparity $E_{\langle i,j\rangle}$ in Eq. 3, after which we detail their construction.
Metric 1: Displacement Vector. To mitigate the effect of noise in raw graphs and features, we estimate the displacement vector $\mathbf{D}_{\langle i,j\rangle}$ using optimized node representations and compute the pairwise distances as
$\mathbf{D}_{\langle i,j\rangle}=\mathbf{h}_{i}-\mathbf{h}_{j},D_{\langle i,j\rangle}=\|\mathbf{h}_{i}-\mathbf{h}_{j}\|_{2},$ (4)
@@
Metric 2: Homophily Ratio Disparity. Considering that recent studies emphasize the importance of capturing the homophily ratio disparity in addressing heterophilous graphs, we weight the neighbor features using this disparity to incorporating both it and distance into the torque formulation. To estimate node-level homophily, it is crucial to annotate the labels of neighboring nodes around a given node. Since labeled data are often scarce, we utilize the model’s final outputs to generate pseudo-labels for unlabeled nodes, with their accuracy improving as the model is progressively optimized. Formally, $E_{\langle i,j\rangle}$ is computed by
$E_{\langle i,j\rangle}=|h_{i}^{+}-h_{j}^{+}|,\ h_{i}^{+}=\frac{|\{v_{j}|\hat{\mathbf{y}}_{i}=\hat{\mathbf{y}}_{j},v_{j}\in\mathcal{N}_{i}\}|}{|\mathcal{N}_{i}|}.$ (5)
Here, $\hat{\mathbf{y}}_{i}$ denotes the ground-truth label for labeled nodes or the pseudo-label for unlabeled nodes. Finally, the torque value of edge $e_{k}$ is computed by:
$T_{e_{k}}=\|\mathbf{D}_{\langle i,j\rangle}\times(E_{\langle i,j\rangle}\mathbf{h}_{j})\|_{2}=\sqrt{D_{\langle i,j\rangle}^{2}\cdot(E_{\langle i,j\rangle}\|\mathbf{h}_{j}\|_{2})^{2}-(E_{\langle i,j\rangle}\mathbf{D}_{\langle i,j\rangle}\cdot\mathbf{h}_{j})^{2}}$ (6)
This formulation captures the combined effects of distance and disparity, facilitating a physics-inspired approach to graph rewiring.
@@ 4.2 Adjust Message Passing
Edge-removal High-order Rewiring. Herein, we propose an automated threshold learning mechanism that identifies the optimal number of edges to prune by pinpointing the largest successive torque gap. Specifically, we first rank all $K$ edges in descending order of their torque values to form a torque‐sorted list (TSL), denoting its $k$-th entry as $\widetilde{e}_{k}$ with torque $\widetilde{T}_{e_{k}}$, so that $\widetilde{T}_{e_{1}}\geq\widetilde{T}_{e_{2}}\geq\cdots\geq\widetilde{T}_{e_{K}}$. We then calculate the torque gap between two consecutive links by
@@
where $\delta$ is a small constant to prevent division by zero, and the weight $\mu_{k}$ reflects the proportion of anomalous edges, those whose distance $D$, disparity $E$ and torque $T$ all exceed their respective means, that are captured within the top $k$ torque-ranked set, emphasizing the boundary between desirable and undesirable connections. The computation formula of $\mu_{k}$ is defined as
@@
where $\bar{D},\bar{E},\bar{T}$ denote the mean values of distance, disparity and torque, respectively, computed over all $K$ edges. The set High_eHigh\_e comprises edges exhibiting above-average values across all three metrics, while Top_kTop\_k contains the top $k$ connections in TSL. According to Eq. 7, we can identify the optimal cutoff by locating the largest torque gap $\mathcal{K}=\arg\max\limits_{0\leq k\leq K-1}G_{k,k+1}$, which separates the edge set into two groups: undesirable connections ($\widetilde{e}_{1},\cdots,\widetilde{e}_{\mathcal{K}}$) and desirable connections ($\widetilde{e}_{\mathcal{K}+1},\cdots,\widetilde{e}_{K}$).
In practice, multi-layer GNNs, such as APPNP [43] and GCNII [27], are widely adopted to enlarge the receptive field of graph convolutions. To allow each layer to adapt the adjacency relationships based on the current information and focus on different structural features of the graph, we adopt a hierarchical rewiring strategy. Building on the pairwise torque gap formulation introduced above, we extend this mechanism across multiple propagation layers. In specific, for each layer $l$, we construct a dedicated propagation matrix that enables selective filtering of undesirable high-order interactions. Let $\mathcal{A}^{(l)}$ denote the refined propagation matrix used in $l$-th layer and $\mathcal{A}^{(0)}=\mathbf{A}$. We identify the current neighbors $v_{j}$ of a given node $v_{i}$ based on $\mathcal{A}^{(l)}$, and recompute the torque as follows
$\mathbf{T}_{\langle i,j\rangle}^{(l+1)}=(\mathbf{h}_{i}^{(l)}-\mathbf{h}_{j}^{(l)})\times E_{\langle i,j\rangle}\mathbf{h}_{j}^{(l)},$ (9)
where $l=0,\cdots,L-1$. Consequently, we gain the (l+1)(l+1)‑th order torque $T^{(l+1)}$ and the corresponding gap $G^{(l+1)}$ using Eqs. 6-8, from which we derive a pruned propagation matrix $\mathcal{A}^{(l+1)^{*}}$ with $(K-\mathcal{K})$ non-zero elements.
Edge-addition High-order Rewiring. In the previous steps, we remove undesirable neighbors by computing the torque of existing edges based on two key attributes. Extending this strategy, we also consider expanding the receptive field by adding edges that are initially absent but potentially beneficial for message passing. However, evaluating torque across all missing edges is computationally intractable, so that we construct a candidate set $\mathcal{T}$ by selecting, for each node, its top-$t$ most similar peers. We then compute the $(l+1)$-th order torque $T^{(l+1)}$ for the resulting $N\times t$ candidate edges, and select $r\times N\times t$ edges with the lowest torque values, where $r$ is a sampling ratio. Nevertheless, this hard selection process is inherently non-differentiable and thus cannot be used in gradient-based optimization. To overcome this, we adopt the Gumbel-Softmax reparameterization trick [44], which enables differentiable sampling by approximating discrete decisions with a continuous relaxation. For each candidate edge $k$, we define its logits $\bm{\pi}_{k}=[\pi_{k0},\pi_{k1}]$, where $\pi_{k0}=T_{e_{k}}^{(l+1)}$ (discard) and $\pi_{k1}=1-T_{e_{k}}^{(l+1)}$ (select). Drawing independent noise $g_{kj}\sim\text{Gumbel (0, 1)}$, the soft selection probabilities are computed via
@@
Messaging Passing on Rewired Graph. To avoid misleading representations in the early stages of training, which could either discard important neighbor information or propagate erroneous signals, rewiring at each layer is always performed with respect to the original input graph. By rewiring the adjacency matrix $\mathbf{A}$ as described, each propagation layer is endowed with an expanded receptive field capable, enabling the capture of effective multi-level interactions.
To evaluate the effectiveness of the proposed THR in capturing high-order information in multi-layer GNNs, we use the deep-based, APPNP, as an example. Subsequent ablation studies and parameter analyses are conducted within this framework. Let $\mathcal{N}_{i}^{(l+1)}=\{v_{j}|\mathcal{A}_{\langle i,j\rangle}^{(l+1)}\neq 0\}$ denotes the refined $(l+1)$-layer neighborhood of node $v_{i}$; then the forward propagation at the $(l+1)$-th layer of APPNP can be reformulated as: